OpenStudy (anonymous):

2x+3y+z=-5 x+2y-z=-6 3x+y+2z=4

OpenStudy (anonymous):

I am not god at all I have failed this 3 times not good

OpenStudy (anonymous):

good

OpenStudy (a_clan):

can be solved by a number of methods . Starting with elimination method. Multiply 2nd eqn with 2 and subtract from 1st. You'll get an equation with 2 variables. Repeat the process with any other pair of equations. Then solve simultaneously.

OpenStudy (anonymous):

do u always mult 2nd equation by 2

OpenStudy (a_clan):

no

OpenStudy (a_clan):

you have to choose a pair of equations and then determine which constant(number) to multiply, so that one variable will cancel out in elimination

OpenStudy (anonymous):

ok i will try that thanks

OpenStudy (amistre64):

x = 1.5 y = -3.1 z = 1.29999.... if i used my ti83 right :)

OpenStudy (anonymous):

i am still lost i am not good at tis at all thx i have to do every step the book does give the ans but says no solution

OpenStudy (amistre64):

every step eh; there are a few ways to solve this, which one are you trying?

OpenStudy (anonymous):

elim but what ever will be the best

OpenStudy (amistre64):

2x +3y + z = -5 x +2y - z = -6 3x + y +2z = 4 Notice that these 2 equaions can eliminate a z 2x +3y + z = -5 x +2y - z = -6 ----------------- 3x +5y = -11 so lets continue eliminating zs :) x +2y - z = -6 ; multiply this one by 2 to get 2 ---------------- 2x+4y -2z = -12 ; now add together 3x + y +2z = 4 ----------------- 5x +5y = -8 hows that so far?

OpenStudy (anonymous):

ok that is great

OpenStudy (anonymous):

the 2 is a neg x i typed it wrong

OpenStudy (amistre64):

this leaves us with 2 new equations to work on 3x +5y = -11 ;there is a 5y in both so lets eliminate those 5x +5y = - 8 ; by muliplying one of them by (-1) 3x +5y = -11 -1 ------------- -3x -5y = 11 5x +5y = - 8 ; now add em :) ------------- 2x = 3 ; and solve for x x = 3/2 = 1.5

OpenStudy (anonymous):

-x+2y-z sorry

OpenStudy (amistre64):

the steps are the same regardess of the numbers; just correct for it :)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

when we correct for that little typo we get: x +5y = -11 x +5y = - 8 these are the same equations but with different paths; in other words they are parallel lines the will never meet. When lines dont meet there cant be a point they have in common for a solution :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u r really good at this

OpenStudy (amistre64):

I can be :) at other times i really mess it up lol

OpenStudy (anonymous):

i have a midterm wed and i know i probally will not pass it

OpenStudy (anonymous):

this is my 3rd time

OpenStudy (anonymous):

i am taking it online and it is harder than class

OpenStudy (amistre64):

college algebra was a breeze for me. oh yeah, online stuff is the pitts

OpenStudy (anonymous):

now i am in big trouble

OpenStudy (anonymous):

we take a test at the school so i cant have my notes

OpenStudy (anonymous):

64m^3x-8n^3x=